OpenStudy (madalton77):

A budget calls for $200 less than one third of the clothing portion of the budget to be spent on the man of the house. If the man needs an estimated $600 for the year how much should be budget for the entire family for the year?

OpenStudy (danjs):

Spent on House = one third clothing spent - 200

OpenStudy (danjs):

Say C = clothing portion H = House portion \[H = \frac{ 1 }{ 3 }C-200\]

OpenStudy (danjs):

right?

OpenStudy (madalton77):

Answer $1800 for the entire family? Correct?

OpenStudy (madalton77):

Less than implies $00 "to the right". Im the mom trying to explain this to my 8th grader.

OpenStudy (madalton77):

I meant less than implies $200 to the right of the equation.

OpenStudy (danjs):

"200 less than 1/3 of the clothing portion" "less than" means 200 fewer, subtract "of" means multiply 1/3 of clothing portion --- 1/3 * C 200 less than that (1/3)*C - 200

OpenStudy (danjs):

\[H = \frac{ 1 }{ 3 }C-200\] So if they say the house needs 600, then you put that value in for H \[600=\frac{ 1 }{ 3 }C-200\] solving for C gives you the clothing budget 800 = 1/3 C C = 2400

OpenStudy (madalton77):

I understand, thank you very much. Its been a while since I have worked on Algebra!!

OpenStudy (danjs):

welcome